The genome of virus SV40 is a circular double‐stranded DNA molecule 5 kb long
(i.e., it contains 10,000 bases

If the number of thymine in the SV40 genome is 1830, what is the percentage
of cytosine in the genome?
1
Expert's answer
2014-05-22T05:54:38-0400
T binds complementary to A, therefore thenumber of T+A is 1830 x 2 = 3 660.
C binds complementary to G, therefore thenumber of C+G is 10 000 - 3 660 = 6 340.
The number of C bases is 6 340 / 2 = 3 170 or 31.7%.